OpenStudy (alina123):

your baby cousin was born at a height of 1 5/12 ft. over the next 5 years she grew to be a height of 3 1/2 ft. find the average yearly growth

ganeshie8 (ganeshie8):

basically u need to find out how much she is growing each year... on average

ganeshie8 (ganeshie8):

to grow from 1 5/12 ft to 3 1/2 ft, she took 5 years eh ?

OpenStudy (alina123):

i would divide by 5

ganeshie8 (ganeshie8):

Exactly ! denominator is 5. what u wil put on the numerator ?

OpenStudy (alina123):

3.5

ganeshie8 (ganeshie8):

think a bit

ganeshie8 (ganeshie8):

actuallt, you need to take the difference

ganeshie8 (ganeshie8):

3.5 - 1 5/12

OpenStudy (alina123):

25 over 12

ganeshie8 (ganeshie8):

Average yearly growth = \(\large \frac{3.5 - 1 \frac{5}{12}}{5}\)

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

\(\large \frac{3.5 - 1 \frac{5}{12}}{5}\) \(\large \frac{3.5 - \frac{17}{12}}{5}\)

ganeshie8 (ganeshie8):

can u simplify further ?

OpenStudy (alina123):

5 over 12

ganeshie8 (ganeshie8):

Correct !! good job :)

OpenStudy (alina123):

thanks bye

ganeshie8 (ganeshie8):

that means, she is growing 5/12 ft per year !

ganeshie8 (ganeshie8):

yw